# Towards Understanding the Fragility of Multilingual LLMs against Fine-Tuning Attacks

### v1

@@ 1 Introduction
Large language models (LLMs) have revolutionized the field of artificial intelligence, but their widespread global adoption has also raised concerns about their safety. Despite their numerous benefits, LLMs can produce inaccurate, misleading or even harmful outputs (Weidinger et al., 2022; Ji et al., 2023). The safety alignment (Ouyang et al., 2022; Wei et al., 2022; Rafailov et al., 2023) of LLMs aims to address safety issues by aligning LLMs to produce outputs that are safe, trustworthy and aligned with human values. However, recent studies have demonstrated that the safety-aligned LLMs is not adversarially robust (Zou et al., 2023; Ghanim et al., 2024; Carlini et al., 2024). In a seminal work, Qi et al. (2023) proposed a fine-tuning attack showing the safety alignment of LLMs can be compromised by fine-tuning only a few steps on a few adversarially designed training examples, either for closed/open-source models (Touvron et al., 2023; Achiam et al., 2023). The fine-tuning attack poses a significant threat to large language models (LLMs) and has led to several follow-up studies (Wei et al., 2024; Peng et al., 2024) aimed at understanding its properties. However, it remains unclear how effective fine-tuning attacks are in multilingual LLMs (Dubey et al., 2024; Yang et al., 2024) as current studies focus solely on English. Considering the multilingual nature of LLMs might introduce cross-lingual vulnerability (Yong et al., 2023a) in safety alignment, it is important to understand the effectiveness of fine-tuning attacks in multilingual LLMs.
@@
To better understand why cross-lingual generalization of fine-tuning attacks exists, we hypothesize that the safety information in safety-aligned multilingual LLMs is language-agnostic. To validate our hypothesis, we propose the method Safety Information Localization (SIL) to localize multilingual safety-related parameters. Our method is inspired by recent work on task knowledge localization (Dai et al., 2022; Panigrahi et al., 2023; He et al., 2024b)—here, we estimate task-specific neuron importance in a manner akin to neuron-pruning (Wei et al., 2024) and Integrated Gradients (Sundararajan et al., 2017). With SIL, we find safety-related information is sparse and shared among different languages—modifying only 20% of an LLM’s weights using monolingual fine-tuning attacks is sufficient to break safety alignment across all languages.
@@ 3 Localizing Language-Agnostic Safety Information
In Section 3, we provide an explanation for the cross-lingual generalization of fine-tuning attacks as observed in Section 2.2. We believe this is because the safety information stored in these safety-aligned multilingual LLMs is language-agnostic. Motivated by recent work that localizes task-specific skills in large models (Dai et al., 2022; Panigrahi et al., 2023; He et al., 2024b), we propose a new localization technique SIL and successfully identify the parameters in these LLMs related to safety knowledge.
@@ 3.2 Stitching with $\bm{\gamma}_{{\text{SIL-}k}}$
We introduce the stitching operation, which uses the binary mask $\bm{\gamma}_{{\text{SIL-}k}}$ to make the safety-aligned pretrained model unsafe: we stitch the selected parameters from the fine-tuned model back onto the pretrained LLM and create grafted LLM, a terminology consistent with previous localization work  (Panigrahi et al., 2023; He et al., 2024b). Here, our goal is to show that stitching $\bm{\gamma}_{{\text{SIL-}k}}$ creates unsafe grafted LLMs. Formally, we refer to the grafted LLM as $\bm{\theta}_{l_{\text{ft}}}^{\text{SIL-}k}$ as shown in Equation 1, where we use $\bm{\gamma}_{{\text{SIL-}k}}$ to stitch the parameters from fine-tuned model $\bm{\theta}_{{l_{\text{ft}}}}$ back to pretrained model $\bm{\theta}_{\text{{pre}}}$. Note that $k$ controls the sparsity of $\bm{\gamma}_{{\text{SIL-}k}}$; the larger the $k$, the more weights in $\bm{\theta}_{\text{{pre}}}$ being changed.
@@
To verify that SIL successfully isolates the safety-related parameters modified by the fine-tuning attack, we compute the violation rate for the grafted LLM, and we compare our results against stitching with parameters localized by two other baselines: Weight-Diff-$k$ and SNIP (Figure 2).
@@
SNIP localization is presented by Wei et al. (2024) to identify safety-critical parameters. We believe SNIP is a special case of SIL, where $\bm{\theta}_{{l_{\text{ft}}}}$ is set to 00. The importance score of each weight in the model is computed as:
@@
Similar to SIL, after localization with SNIP, we binarize the result selecting the top-$k$ importance score to be set to $1$ in the binary mask $\bm{\gamma}_{\text{{SNIP-{$k$}}}}$.
@@ Results
SIL is a superior localization technique compared to Weight-Diff-$k$ and SNIP, as Figure 2 shows that we need less parameters to stitch in order to make the pretrained models exhibit high violation rate. One reason is that SIL leverages the gradient information, which proves vital in mitigating the task interference observed in the Weight-Diff-$k$ approach (Panigrahi et al., 2023). Another reason is that SIL considers the influence of parameters shift from the safety-aligned $\bm{\theta}_{\text{{pre}}}$ to $\bm{\theta}_{{l_{\text{ft}}}}$, whereas SNIP misses this crucial information of a specific fine-tuned models. Due to the advantages of SIL over other baselines, we use it as the localization method in the following experiments.
@@
Figure 2: Violation rate vs. sparsity $k$ with SIL, SNIP, and Weight-Diff-$k$ methods, for Qwen-2-7B (top) and Llama-3.1-8B (bottom). When choosing $k=20\%$, SIL have the similar VR to the fine-tuned models.
@@ 3.3 Is the safety information stored in the model language-agnostic?
In this subsection we understand whether the safety information stored in the model is language-agnostic. We leverage the localized parameters to give insights to why fine-tuning in one language can disrupt the safety of all languages. We hypothesize that, if different mask vectors (say $\bm{\gamma}_{{l_{0}}}$ and $\bm{\gamma}_{{l_{1}}}$) share similar parameters, then the information represented by these parameters is likely important across all such masks, thereby reducing dependency on specific languages, like $l_{0}$ and $l_{1}$. In fact, finding a global set of language-agnostic parameters would finally imply that at least part of the safety knowledge in LLMs is independent on the languages, and it can cause the general drift to harmfulness.
@@
If multilingual LLMs do encode language-agnostic knowledge about safety, then the shared safety information between two languages (i.e., $\text{SIR}_{l_{0},l_{1}}$) must be large. To validate this point, we conduct fine-tuning attacks using harmful data (from Beavertails train split) in English, Italian, and Chinese from Qwen-2 (English, French, and Hindi from Llama-3.1), and compute SIL-20 masks using calibration data (from Beavertails test split) in five languages. Then, we compute the bilingual SIR between $3\times 5$ times (three languages used to fine-tune the models plus two additional languages).
@@
We show that the bilingual SIR value between the masks obtained from the harmful calibration data is substantially larger than the benign (Table 1) and random baselines (which settles at 20% by construction). It is also worth pointing out the bilingual SIR computed with the benign baseline in each row in Table 1 shares the same language used to fine-tuned the model. The result suggests that fine-tuning attacks in one language impact the safety-related parameters of different languages, more than they do to other types of parameters (even for the helpfulness-related parameters in the same languages). Figures 3 and 6 further validate these findings: stitching the bilingual intersections of localized parameters (e.g.$\bm{\gamma}_{\text{{EN}}}\cap\bm{\gamma}_{\text{{IT}}}$) back onto the original safety-aligned multilingual LLMs (e.g.$\bm{\theta}_{\text{EN}}^{\text{EN}\cap\text{IT}}$) (orange bars) reports similarly large violation rates as the jailbroken fine-tuned models $\bm{\theta}_{{l_{\text{ft}}}}$ (blue bar), whereas the benign baseline $\bm{\theta}_{{\text{Benign}_{l_{\text{ft}}}}}$ (green bar) and the original safety-aligned multilingual LLMs $\bm{\theta}_{\text{{pre}}}$ (red bar) remain safe. Moreover, we hypothesize that the preference for the English language showed in Table 1 by Llama-3.1-8B, can be explained by the findings in Wendler et al. (2024), where it is demonstrated that the “concept space” in the models of the Llama family is more closely aligned with English than with other languages (Table 2 also suggests similar results).
@@
Table 2: Multilingual (global) SIR results. Even removing a massive amount of language-dependent knowledge, SIL localized parameters share more language-agnostic safety information than when compared to the benign and the random baselines.
@@ 4.1 Explanation for why freezing safety-related parameters fails to prevent fine-tuning attacks
Table 3: Multilingual (global) SIR results after parameter freezing (indicated by overlines over the metrics). The new language-agnostic parameters has zero intersection with the one obtained without freezing during fine-tuning. Again, it shows to share a very large volume of safety information, when compared to the benign and random baselines.
@@ 5 Related Work
LLM safety alignment through instruction-tuning and RLHF (Wei et al., 2021; Ouyang et al., 2022; Touvron et al., 2023) aims to align the behaviors of LLMs align with human values, Jailbreaking a safety-aligned model aims at bypassing or removing these safety guardrails. It can be realized either by only modifying the prompts (Liu et al., 2023a, b; Zou et al., 2023), or further fine-tuning (Qi et al., 2023; Zhan et al., 2023; Poppi et al., 2024). In terms of fine-tuning attacks, Peng et al. (2024) study fine-tuning attacks by randomly perturbing model weight parameters and finds out that safety alignment of LLMs is easily broken if the the model weights are away from the “safety basin” in parameter weight space. He et al. (2024a) select benign data strategically for fine-tuning attacks.
@@
The model parameter space offers a simple perspective for task localization and knowledge attribution, as it represents the landscape of all possible models with a given structure. A variety of different works observed the models’ tendency of mapping knowledge into specific parameters in the model parameter space (Bereska and Gavves, 2024). In particular, Hao et al. (2021) and Dai et al. (2022) leverage Integrated Gradients proposed in Sundararajan et al. (2017) originally used for input feature attribution, and modify it accordingly to analyze relational facts. Wei et al. (2024) reuse of neuron pruning proposed by Lee et al. (2019) to localize safety relevant weights, and show removing these weight from the pre-trained model pushes it back to an unsafe status. Inspired by those methods, we identify language-agnostic safety parameters in the model parameter space by estimating language-specific neuron importance in a manner akin to neuron-pruning (Wei et al., 2024) and Integrated Gradients (Sundararajan et al., 2017).
@@
Multilingual LLMs and its safety problem are gaining increasing attention. Unlike detoxification (Li et al., 2024), safety refusal has poor cross-lingual generalization. For instance, simply translating English malicious prompts into non-English can bypass safety guardrails in both closed-source and open-source LLMs (Yong et al., 2023a; Wang et al., 2023; Deng et al., 2023). Linguistic form changes, such as transliteration (Ghanim et al., 2024) and code-switching (Upadhayay and Behzadan, 2024), can also enable jailbreaking of safety guardrails. Furthermore, Shen et al. (2024) demonstrate that English refusal training generalizes poorly for both low-resource and high-resource languages such as Mandarin Chinese. Our research further demonstrates the cross-lingual fragility of safety refusal guardrails, illustrating that fine-tuning attacks in one language can compromise LLMs across multiple languages due to the language-agnostic safety knowledge embedded in safety-aligned LLMs.

### v2

@@ 1 Introduction
Large language models (LLMs) have revolutionized the field of artificial intelligence, but their widespread global adoption has also raised concerns about their safety. Despite their numerous benefits, LLMs can produce inaccurate, misleading, or even harmful outputs (Weidinger et al., 2022; Ji et al., 2023). The safety alignment (Ouyang et al., 2022; Wei et al., 2022; Rafailov et al., 2023) of LLMs aims to address safety issues by aligning LLMs to produce outputs that are safe, trustworthy and aligned with human values. However, recent studies have demonstrated that the safety-aligned LLMs are not adversarially robust (Zou et al., 2023; Ghanim et al., 2024; Carlini et al., 2024). In a seminal work, Qi et al. (2023) proposed a fine-tuning attack showing the safety alignment of LLMs can be compromised by fine-tuning only a few steps on a few adversarially designed training examples, either for closed/open-source models (Touvron et al., 2023; Achiam et al., 2023). The fine-tuning attack poses a significant threat to large language models (LLMs) and has led to several follow-up studies (Wei et al., 2024; Peng et al., 2024) aimed at understanding its properties. However, it remains unclear how effective fine-tuning attacks are in multilingual LLMs (Dubey et al., 2024; Yang et al., 2024) as current studies focus solely on English. Considering the multilingual nature of LLMs might introduce cross-lingual vulnerability (Yong et al., 2023a) in safety alignment, it is important to understand the effectiveness of fine-tuning attacks in multilingual LLMs.
@@
To better understand why cross-lingual generalization of fine-tuning attacks exists, we hypothesize that the safety information in safety-aligned multilingual LLMs is language-agnostic. To validate our hypothesis, we propose the method Safety Information Localization (SIL) to localize multilingual safety-related parameters. Our method is inspired by recent work on task knowledge localization (Dai et al., 2022; Panigrahi et al., 2023; He et al., 2024c)—here, we estimate task-specific neuron importance in a manner akin to neuron-pruning (Wei et al., 2024) and Integrated Gradients (Sundararajan et al., 2017). With SIL, we find safety-related information is sparse and shared among different languages—modifying only 20% of an LLM’s weights using monolingual fine-tuning attacks is sufficient to break safety alignment across all languages.
@@ 3 Localizing Language-Agnostic Safety Information
In Section 3, we provide an explanation for the cross-lingual generalization of fine-tuning attacks as observed in Section 2.2. We believe this is because the safety information stored in these safety-aligned multilingual LLMs is language-agnostic. Motivated by recent work that localizes task-specific skills in large models (Dai et al., 2022; Panigrahi et al., 2023; He et al., 2024c), we propose a new localization technique SIL and successfully identify the parameters in these LLMs related to safety knowledge.
@@ 3.2 Stitching with $\bm{\gamma}_{{\text{SIL-}k}}$
We introduce the stitching operation, which uses the binary mask $\bm{\gamma}_{{\text{SIL-}k}}$ to make the safety-aligned pretrained model unsafe: we stitch the selected parameters from the fine-tuned model back onto the pretrained LLM and create grafted LLM, a terminology consistent with previous localization work  (Panigrahi et al., 2023; He et al., 2024c). Here, our goal is to show that stitching $\bm{\gamma}_{{\text{SIL-}k}}$ creates unsafe grafted LLMs. Formally, we refer to the grafted LLM as $\bm{\theta}_{l_{\text{ft}}}^{\text{SIL-}k}$ as shown in Equation 1, where we use $\bm{\gamma}_{{\text{SIL-}k}}$ to stitch the parameters from fine-tuned model $\bm{\theta}_{{l_{\text{ft}}}}$ back to pretrained model $\bm{\theta}_{\text{{pre}}}$. Note that $k$ controls the sparsity of $\bm{\gamma}_{{\text{SIL-}k}}$; the larger the $k$, the more weights in $\bm{\theta}_{\text{{pre}}}$ being changed.
@@
To verify that SIL successfully isolates the safety-related parameters modified by the fine-tuning attack, we compute the violation rate for the grafted LLM, and compare our results against stitching with parameters localized by two other baselines: Weight-Diff-$k$ and SNIP (Figure 2).
@@
SNIP localization is presented by Wei et al. (2024) to identify safety-critical parameters. We believe that SNIP is a special case of SIL, where $\bm{\theta}_{{l_{\text{ft}}}}$ is set to 00. The importance score of each weight in the model is computed as:
@@
Similarly to SIL, after localization with SNIP, we binarize the result selecting the top-$k$ importance score to be set to $1$ in the binary mask $\bm{\gamma}_{\text{{SNIP-{$k$}}}}$.
@@ Results
SIL is a superior localization technique compared to Weight-Diff-$k$ and SNIP, as Figure 2 shows that we need less parameters to stitch in order to make the pretrained models exhibit high violation rate. One reason is that SIL leverages the gradient information, which is proved vital in mitigating the task interference observed in the Weight-Diff-$k$ approach (Panigrahi et al., 2023). Another reason is that SIL considers the influence of parameters shift from the safety-aligned $\bm{\theta}_{\text{{pre}}}$ to $\bm{\theta}_{{l_{\text{ft}}}}$, whereas SNIP misses this crucial information of a specific fine-tuned models. Due to the advantages of SIL over other baselines, we use it as the localization method in the following experiments.
@@
Figure 2: Violation rate vs. sparsity $k$ with SIL, SNIP, and Weight-Diff-$k$ methods, for Qwen-2-7B (left) and Llama-3.1-8B (right). When choosing $k=20\%$, SIL have the similar VR to the fine-tuned models.
@@ 3.3 Is the safety information stored in the model language-agnostic?
In this subsection we understand whether the safety information stored in the model is language-agnostic. We leverage the localized parameters to give insights into why fine-tuning in one language can disrupt the safety of all languages. We hypothesize that, if different mask vectors (say $\bm{\gamma}_{{l_{0}}}$ and $\bm{\gamma}_{{l_{1}}}$) share similar parameters, then the information represented by these parameters is likely important across all such masks, thereby reducing dependency on specific languages, like $l_{0}$ and $l_{1}$. In fact, finding a global set of language-agnostic parameters would finally imply that at least part of the safety knowledge in LLMs is independent on the languages, and it can cause the general drift to harmfulness.
@@
If multilingual LLMs encode language-agnostic knowledge about safety, then the shared safety information between two languages (i.e., $\text{SIR}_{l_{0},l_{1}}$) must be large. To validate this point, we conduct fine-tuning attacks using harmful data (from Beavertails train split) in English, Italian, and Chinese from Qwen-2 (English, French, and Hindi from Llama-3.1), and compute SIL-20 masks using calibration data (from Beavertails test split) in five languages. Then, we compute the bilingual SIR between $3\times 5$ times (three languages used to fine-tune the models plus two additional languages).
@@
We show that the bilingual SIR value between the masks obtained from the harmful calibration data is substantially larger than the benign (Table 1) and random baselines (which settles at 20% by construction). It is also worth pointing out the bilingual SIR computed with the benign baseline in each row in Table 1 shares the same language used to fine-tuned the model. The result suggests that fine-tuning attacks in one language impact the safety-related parameters of different languages, more than they do to other types of parameters (even for the helpfulness-related parameters in the same languages).
Figures 3 and 6 further validate these findings: stitching the bilingual intersections of localized parameters $\bm{\gamma}_{\text{{EN}}}\cap\bm{\gamma}_{\text{{IT}}}$ back onto the original safety-aligned multilingual LLMs $\bm{\theta}_{\text{EN}}^{\text{EN}\cap\text{IT}}$ (orange bars) reports similarly large violation rates as the jailbroken fine-tuned models $\bm{\theta}_{{l_{\text{ft}}}}$ (blue bar), whereas the benign baseline $\bm{\theta}_{{\text{Benign}_{l_{\text{ft}}}}}$ (green bar) and the original safety-aligned multilingual LLMs $\bm{\theta}_{\text{{pre}}}$ (red bar) remain safe. Moreover, we hypothesize that the preference for the English language showed in Table 1 by Llama-3.1-8B, can be explained by the findings in Wendler et al. (2024), where it is demonstrated that the “concept space” in the models of the Llama family is more closely aligned with English than with other languages (Table 2 also suggests similar results).
We further analyze the relationship between the bilingual SIR and the violation rate observed across languages. In particular, we observe that, despite the bilingual SIR overlap between Chinese and English (69.7% in Qwen-2) is lower than the overlap between Chinese and itself (100% in Qwen-2), the violation rate of the model fine-tuned in Chinese when tested in English is higher than when tested in Chinese (Figure 1). This suggests that while many safety-related parameters are shared across languages, their actual influence on model behavior may vary. Specifically, fine-tuning harmful data in Chinese may have localized effects that preserve more of the original safety constraints, whereas English may be more susceptible to degradation. Moreover, additional factors can exacerbate the discrepancy between SIR and violation rate: First, the harmfulness detector sensitivity to different languages may influence the reported violation rates; Second, linguistic characteristics, such as sentence structure, vary significantly between different languages, thus affecting how well the safety capabilities generalize from one language to another.
@@
Table 2: Multilingual (global) SIR results. Even removing a massive amount of language-dependent knowledge, SIL localized parameters share more language-agnostic safety information than when compared to the benign baselines.
@@ 4.1 Explanation for why freezing safety-related parameters fails to prevent fine-tuning attacks
Table 3: Multilingual (global) SIR results after parameter freezing (indicated by overlines over the metrics). The new language-agnostic parameters has zero intersection with the one obtained without freezing during fine-tuning. Again, it shows to share a very large volume of safety information, when compared to the benign baselines.
@@ 5 Related Work
LLM safety alignment through instruction-tuning and RLHF (Wei et al., 2021; Ouyang et al., 2022; Touvron et al., 2023) aims to align the behaviors of LLMs with human values. Jailbreaking a safety-aligned model aims at bypassing or removing these safety guardrails. It can be achieved either by only modifying the prompts (Liu et al., 2023a, b; Zou et al., 2023), or further fine-tuning (Qi et al., 2023; Zhan et al., 2023; Poppi et al., 2024).
In terms of fine-tuning attacks, Peng et al. (2024) study fine-tuning attacks by randomly perturbing model weight parameters and find that safety alignment of LLMs is easily broken if the model weights deviate from the “safety basin” in parameter weight space. He et al. (2024a) strategically select benign data for fine-tuning attacks. In contrast, our work focuses on identifying safety-relevant parameters and analyzing the impact of multilingual fine-tuning attacks from a mechanistic perspective.
@@
The model parameter space offers a fundamental perspective for task localization and knowledge attribution, as it represents the landscape of all possible models with a given structure. A variety of studies have observed models’ tendency to encode specific knowledge into distinct parameters in the parameter space (Bereska and Gavves, 2024). In particular, Hao et al. (2021) and Dai et al. (2022) leverage Integrated Gradients (Sundararajan et al., 2017), originally used for input feature attribution, and modify it to analyze relational facts. Wei et al. (2024) reuse neuron pruning (Lee et al., 2019) to identify safety-relevant parameters, demonstrating that removing these parameters pushes a pre-trained model back to an unsafe state. Arditi et al. (2024) also study safety mechanisms in LLMs, they focus on representation space rather than parameter space, which is the primary concern of our work. Their approach identifies critical directions in the activation space rather than pinpointing where in the LLM safety-related parameters reside. This fundamental distinction allows our method to directly analyze and manipulate the parameters responsible for safety alignment. Additionally, their study does not address multilingual safety, whereas we focus on cross-lingual safety alignment.
Inspired by these prior approaches, our work identifies language-agnostic safety parameters in the model parameter space by estimating language-specific neuron importance, akin to neuron pruning (Wei et al., 2024) and Integrated Gradients (Sundararajan et al., 2017). Through this approach, we provide a mechanistic explanation for cross-lingual vulnerabilities in safety alignment.
@@
The safety of multilingual LLMs is a growing area of concern. Unlike detoxification approaches (Li et al., 2024), safety refusal exhibits poor cross-lingual generalization. Translating English adversarial prompts into non-English languages can often bypass safety guardrails in both proprietary and open-source models (Yong et al., 2023a; Wang et al., 2023; Deng et al., 2023). Other linguistic transformations, such as transliteration (Ghanim et al., 2024) and code-switching (Upadhayay and Behzadan, 2024), further enable jailbreaking of safety mechanisms.
Furthermore, Shen et al. (2024) show that English safety refusal training generalizes poorly, even for high-resource languages such as Mandarin Chinese. Our work extends these findings by demonstrating that fine-tuning attacks in one language can compromise safety alignment across multiple languages due to the shared, language-agnostic nature of safety-related parameters in multilingual LLMs.
One contemporary work also investigates cross-lingual vulnerabilities (He et al., 2024b). While both our work and theirs show that fine-tuning in one language can lead to safety degradation across languages, their study lacks a mechanistic explanation for why this occurs. Our contributions go beyond merely presenting the attack—we further explain cross-lingual generalization using mechanistic interpretability methods and introduce a cross-lingual jailbreak method that attacks LLMs adapted to new languages. While He et al. (2024b) primarily study backdoor attacks by substituting benign fine-tuning datasets with adversarially fabricated responses (e.g., responses containing explicit hate speech triggers), we consider natural-language, multilingual prompts and harmful assistant responses that more closely resemble real-world fine-tuning vulnerabilities and better capture practical adversarial fine-tuning risks. Finally, our method of localizing safety-relevant parameters allows us to confirm the alternative pathways hypothesis (Wei et al., 2024).